### v2

# Louis Pallegoix1, Jaime Travesedo1, Alexandre S. May1,2, Léo Balembois1, Denis Vion1, Patrice Bertet*1 and Emmanuel Flurin Université Paris-Saclay, CEA, CNRS, SPEC, 91191 Gif-sur-Yvette Cedex, France 2Alice&\&Bob, 53 boulevard du Général Martial Valin, 75015 Paris

Abstract
We report on the characteristics of a microwave photon counter device based on a superconducting transmon qubit. Its design is similar to [4], with an additional bandwidth tuning circuit that permits tuning the detection bandwidth over an order of magnitude to optimize the device efficiency and noise. Owing to this new feature and to improvements in device fabrication, a power sensitivity of $3\cdot 10^{-23}\ \mathrm{W}/\sqrt{\mathrm{Hz}}$ is reached. We confirm the high performance of the device by measuring single spin microwave fluorescence.
I Introduction
Single-photon detection has long been a fundamental tool in quantum technologies, with well-established applications in the optical domain, ranging from fluorescence microscopy [32, 22, 9, 14] to quantum communication [18] and boson sampling [5] ; however, extending single-photon detection into the microwave regime presents significant challenges. The energy of microwave photons—of the order of µeV—is several orders of magnitude lower than that of optical photons, requiring detectors to operate at cryogenic temperatures to suppress thermal noise and reduce the background thermal photon population.
Despite these challenges, single-microwave-photon detectors (SMPDs) have emerged as critical tools for advancing quantum sensing and information processing. Their ability to detect individual microwave photons has opened possibilities for a wide range of applications, including the detection of weak incoherent microwave emitters, such as electron spins in solids [1, 41, 10], and even searches for hypothetical dark-matter particles [26, 16]. SMPDs also hold potential for primary thermometry at low temperatures [38] and for use in quantum illumination protocols [3]. Moreover, these detectors can play a key role in quantum information processing [33, 13, 6], enabling heralded entanglement between superconducting qubits [29], improving qubit readout schemes [31], and facilitating the generation of robust quantum states [7].
Various SMPD designs have been proposed and implemented, with most relying on superconducting qubits [28, 1, 29, 34, 20, 36, 25, 37, 19, 35, 15, 24, 21, 8, 23], quantum dots [42, 17] or bolometric detectors [27]. These devices have demonstrated promising results, achieving sensitivities ranging from $10^{-21}$ to $10^{-22}\,\mathrm{W/\sqrt{Hz}}$ at microwave frequencies ; however, there remain limitations in terms of bandwidth, detection efficiency, and dark count rates that must be addressed to further enhance their performance.
[…]
II SMPD principle of operation
II.1 Device design
The SMPD device (see Figs. 1a and b) consists of a transmon qubit (mode $q$, frequency $\omega_{q}$), capacitively coupled to two harmonic oscillator modes, called the buffer (mode $b$, frequency $\omega_{b}$), and the waste (mode $w$, frequency $\omega_{w}$). The buffer is connected to the input line of the device through another resonator at frequency $\omega_{pb}$. The latter serves as Purcell filter to decouple the qubit from the input line, while also setting the effective buffer energy coupling rate $\kappa_{\mathrm{b,c}}$. The total energy-loss rate of the buffer resonator is $\kappa_{b}=\kappa_{\mathrm{b,c}}+\kappa_{\mathrm{b,i}}$, where $\kappa_{\mathrm{b,i}}$ represents the internal losses of the buffer. The waste is connected to an output line through another Purcell filter, at frequency $\omega_{pw}$, with a rate $\kappa_{w}$. Other relevant parameters are the qubit energy-relaxation time $T_{1}$ and equilibrium excited-state population, $p_{\mathrm{th},q}$, which varies with the qubit temperature $T$, in the $T\ll\hbar\omega/k_{B}$ limit, as $p_{\mathrm{th},q}=\frac{1}{e^{\frac{\hbar\omega_{q}}{k_{B}T}}-1}$. In this formula, $T$ is not necessarily the cryostat base temperature, as various heating effects (due to infrared radiation in particular) may lead to a higher effective qubit temperature.
For optimal operation as an energy detector, both the frequency and bandwidth of the device need to be matched to the source of interest, be it a spin [4, 41] or a dark-matter candidate [12]. Frequency tuning is achieved by inserting a superconducting quantum interference device (SQUID) into the buffer mode, resulting in a buffer frequency $\omega_{b}(\phi_{b})$ that depends on the flux threading the buffer SQUID loop, $\phi_{b}$. An asymmetric SQUID is purposely chosen to minimize sensitivity to flux noise. Bandwidth tuning is achieved by inserting a second SQUID into the Purcell filter mode, resulting in a buffer coupling rate $\kappa_{\mathrm{b,c}}(\phi_{\mathrm{pb}})$ that depends on the flux threading the Purcell SQUID loop, $\phi_{\mathrm{pb}}$. This SQUID is symmetric to maximize the bandwidth-tuning range.
II.2 Four-wave mixing
The working principle of photon detection relies on a 4WM process that maps the presence or absence of a photon onto the state of the transmon qubit. The nonlinear element is provided by the qubit Josephson junction. The 4WM is activated by a microwave signal (called the pump, frequency $\omega_{p}$) of amplitude $\xi_{0}$, applied to the qubit via a pump line connected capacitively to the qubit (see Fig. 1). When the pump frequency approximately satisfies the relation $\omega_{b}+\omega_{p}=\omega_{q}+\omega_{w}-\chi_{w}$, the dominant process is described by the Hamiltonian
$H_{\mathrm{4WM}}=\sqrt{\chi_{b}\chi_{w}}(\xi_{0}bq^{\dagger}w^{\dagger}+\xi_{0}^{*}b^{\dagger}qw),$ (1)
where $\chi_{b}$ (resp. $\chi_{w}$) is the qubit dispersive coupling to the buffer mode (resp. the waste mode) [28]. This Hamiltonian reversibly converts a buffer and pump photon into a qubit excitation and waste photon. Dissipation of the waste photon into the output line ensures that the process is irreversible ; nonetheless, the qubit remains excited long enough to be readout subsequently, and it can thus be used as a flag that frequency conversion occurred.
An approximate description is obtained by eliminating the qubit degree of freedom, yielding an effective Hamiltonian $H_{\mathrm{eff}}=g_{\mathrm{4WM}}bw^{\dagger}+h.c.$ with $g_{\mathrm{4WM}}=\sqrt{\chi_{b}\chi_{w}}\xi_{0}$. This describes the parametric coupling between the buffer and waste modes, similar to the Josephson parametric converter for instance. Assuming that $\kappa_{\mathrm{b,i}}\ll\kappa_{\mathrm{b,c}}$, the efficiency of frequency conversion under this process is $\eta_{\mathrm{4WM}}=4C/(1+C)^{2}$, where $C=\frac{4\chi_{b}\chi_{w}|\xi_{0}|^{2}}{\kappa_{w}\kappa_{b}}$ is the system cooperativity. Unit efficiency is reached when $C=1$. Physically, this occurs when the frequency-conversion rate $4|g_{\mathrm{4WM}}|^{2}/\kappa_{w}$ is equal to the input resonator coupling rate, $\kappa_{b}$. This condition can always be achieved by adjusting the pump frequency and amplitude (see Appendix A for more details).
[…]
II.3 Cyclic operation
[…]
In the third step, the qubit is reset into its ground state by applying a $\pi$ pulse conditioned to the outcome of the previous readout being the excited state. This efficiently resets the qubit into its ground state, with a probability that can be even lower than the thermal-equilibrium qubit excited-state probability [4]. Note that the duration of the reset step $T_{\mathrm{reset}}$ varies from cycle to cycle due to its conditional nature. The detector duty cycle $\eta_{\mathrm{cycle}}$ is defined as the average value of $T_{d}/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}})$.
III Figures of merit
In this section, we describe in some detail the processes that limit the performance of the device as an energy detector. This performance is captured by the power sensitivity $S$, defined as the power that can be detected with unit signal-to-noise ratio in $1$ s integration time [4]:
$S=\hbar\omega_{b}\frac{\sqrt{\alpha}}{\eta_{\mathrm{SMPD}}}.$ (2)
In this formula, $\eta_{\mathrm{SMPD}}$ is the overall detection efficiency of the SMPD and $\alpha$ is the average number of counts per second when no signal is sent to the device. The latter can be separated in two contributions $\alpha=\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$, $\alpha_{\mathrm{err}}$ being the false-positive events and $\alpha_{\mathrm{th}}$ being the counts due the detection of thermal photons in the input line. In what follows, we address each of these three quantities, with the goal of maximizing $\eta_{\mathrm{SMPD}}$ while minimizing $\alpha_{\mathrm{err}}$ and $\alpha_{\mathrm{th}}$.
III.1 Detection efficiency
A photon can be missed if :
  • It is spectrally filtered by the detector
  • The 4WM efficiency is not 11
  • The qubit relaxes back in its ground state during the detection step
  • Qubit excited state readout fails
  • The photon hits the detector during the readout or reset steps
Therefore, the probability that a photon in the input line at frequency $\omega$ gives rise to a count is
[…]
where the detector frequency response function is approximated by a Lorentzian of width $\kappa_{d}$ (the detector linewidth) such that $\eta_{\omega}=\frac{1}{1+\left[\frac{2(\omega-\omega_{b})}{\kappa_{d}}\right]^{2}}$ . The qubit efficiency $\eta_{q}=\frac{T_{1}}{T_{d}}(1-\mathrm{e}^{-\frac{T_{d}}{T_{1}}})$ can be approximated by $1-\frac{T_{d}}{2T_{1}}$ in the limit $T_{d}\ll T_{1}$ [4]. To maximize $\eta_{\mathrm{SMPD}}$, $T_{d}$ should be optimized to maximize the duty cycle while keeping the cycle time well below $T_{1}$ to avoid relaxation of the qubit before readout.
III.2 Dark count rate
The dark count rate $\alpha_{\mathrm{err}}$ is the number of counts detected per unit of time, even though no photon was present in the input line. It has two contributions, $\alpha_{\mathrm{err}}=\alpha_{q}+\alpha_{p}$. First, the qubit can be found in the excited state despite the active reset procedure. This is due to the relaxation toward the thermal-equilibrium value $p_{q,th}$, which takes place during the detection window with a characteristic time $T_{1}$ [4], giving rise to $\alpha_{q}=p_{th,q}(T_{d}/T_{1})(1/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}}))$ (assuming that the active reset works perfectly, otherwise an offset must be added to that expression). The other contribution $\alpha_{p}$ is due to heating brought by the pump. Note that the readout errors causing the ground state to be erroneously detected as the excited state are negligible because we purposely choose a detection threshold that minimizes this effect.
III.3 Thermal counts
Thermalizing the microwave field at millikelvin temperatures is well known to be a difficult task because of vanishing thermal conductivity and the large attenuation needed in the input lines. This results in an effective field temperature that is often significantly larger than the cryostat base temperature, and in our case, in the presence of a finite flux of thermal photons at the SMPD input, causing a thermal count rate of thermal origin
$\alpha_{\mathrm{th}}=\frac{1}{2\pi}\int_{-\infty}^{\infty}\bar{n}_{\mathrm{th,b}}\eta_{\mathrm{SMPD}}(\omega)d\omega,$ (4)
where $\bar{n}_{\mathrm{th,b}}=\frac{1}{e^{\frac{\hbar\omega_{b}}{k_{B}T}}-1}$ is the number of photons per mode assuming that the field is at temperature $T$. Because $\kappa_{d}\ll\omega_{b}$, we get
$\alpha_{\mathrm{th}}=\frac{\bar{n}_{\mathrm{th,b}}\kappa_{d}\eta_{\mathrm{SMPD}}(\omega_{b})}{4}.$ (5)
Due to the tunable bandwidth of our device, we are able here to test the predicted proportionality between noise and detector bandwidth, as will be seen below.
IV Measurements
IV.1 Device characterization
The SMPD resonators are measured with a vector network analyser (VNA). Each resonance is fitted to determine the resonator frequency and linewidth. The results of these fits are reported in Tab. 1.
Resonator Frequency Linewidth
Buffer Purcell $\omega_{pb}/2\pi\in[7.26;7.78]$ GHz 20 MHz
Buffer $\omega_{b}/2\pi\in[7.70;7.76]$ GHz See fig. 2
Qubit $\omega_{q}/2\pi=6.533$ GHz 16 kHz
Waste $\omega_{w}/2\pi=8.475$ GHz 1.75 MHz
Waste Purcell $\omega_{pw}/2\pi\approx 8.39$ GHz 400 MHz
Table 1: Measured resonator frequencies and linewidths. The Purcell filter of the buffer and the buffer itself are tunable over the indicated interval. The buffer linewidth depends on the the Purcell frequency, as shown in fig. 2.
Resonator Cross-Kerr to qubit
Buffer $\chi_{b}/2\pi=3.5$ MHz
Waste $\chi_{w}/2\pi=16$ MHz
Table 2: Measured dispersive shifts induced by the qubit to its neighboring resonators.
Figure 2: Tunability curves. a) Solid circles are the measured frequency of the buffer resonator, $\omega_{b}$, as a function of the flux bias applied to its SQUID, $\Phi_{b}$, for several values of $\Phi_{pb}$ (blue is $0.45\Phi_{0}$, orange is $0.24\Phi_{0}$, green is $0.03\Phi_{0}$). The solid lines are sinusoidal fits, yielding the asymmetry between the buffer SQUID junctions. b) Full blue circles are the measured frequency of the Purcell filter, $\omega_{pb}$, as a function of the flux bias sent to its SQUID $\Phi_{pb}$. The solid line is a fit assuming a symmetric SQUID. c) Full circles show the measured coupling rate of the buffer resonator to the $50\Omega$ line, as a function of its detuning from the Purcell filter, $\omega_{b}-\omega_{pb}$, for 33 different values of $\Phi_{b}$ : blue for $\Phi_{b}=0$, orange for $\Phi_{b}=0.25\Phi_{0}$, green for $\Phi_{b}=0.5\Phi_{0}$. The dashed line shows the internal losses of the buffer, $\kappa_{b,i}$. The solid lines are lorentzian fits to data of the same color. We can see that $\kappa_{b}/2\pi$ rise above 3 MHz at zero detuning and fall below $\kappa_{b,i}$ when the detuning exceeds 100 MHz in absolute value. d) Full circles are the measured SMPD efficiency, $\eta_{\mathrm{SMPD}}$, as a function of signal frequency $\omega$, for $\kappa_{b}/2\pi=170$ kHz (blue) and $280$ kHz (orange). Dashed lines are Lorentzian fits yielding the SMPD bandwidth $\kappa_{d}/2\pi$, equal to $314$ kHz and $531$ kHz, respectively.
Figure 2a shows $\omega_{b}(\phi_{b})$ for several values of $\phi_{pb}$. The buffer tuning range is narrow (60 MHz) and the curve is well fitted by a sinusoid [11], from which we deduce a ratio of 15 between the Josephson-junction critical currents of the asymmetric SQUID. Little crosstalk between the $\phi_{b}$ and $\phi_{pb}$ is observed, as evidenced by the weak dependence of $\omega_{b}/2\pi$ on $\phi_{pb}$.
From the fits of the buffer resonance curves, we extract $\kappa_{b,i}=2.2\cdot 10^{5}\mathrm{s}^{-1}$, and $\kappa_{b,c}$, the internal loss and coupling rates, respectively. Fig. 2c shows $\kappa_{b,c}/2\pi$ as a function of the detuning from the Purcell, $\omega_{b}-\omega_{pb}$. A large tuning range is observed, ranging from $\sim 10$ kHz up to $\sim 3$ MHz. The tuning curve is identical for all values of buffer-resonator frequency (see Fig. 2c). The buffer is overcoupled over a large parameter range (corresponding to $\kappa_{b,c}\gg\kappa_{b,i}$), which is desirable for SMPD operation, as buffer internal losses further reduce $\eta_{\mathrm{4WM}}$ [4].
[…]
IV.2 Four-wave-mixing tuning
The 4WM condition is found by sending a pump tone at frequency $\omega_{p}$ together with a signal at frequency $\omega$ into the SMPD input, followed by qubit readout. Sweeping $\omega_{p}$ and $\omega$, the 4WM condition is identified by an enhanced qubit excited state when the signal is present, which vanishes when it is switched off (see Fig. 7) ; the pump power is then tuned to reach unit cooperativity (see Appendix and Fig. 7).
IV.3 Cyclic operation : Detector bandwidth and dark counts
We now turn to the operation of the SMPD in cyclic mode. The detection step duration is set to $T_{d}=15$ µs. When no photon is sent to the device, the qubit is most often in the ground state at the end of the sequence, implying that no reset is needed, and that the average cycle duration is very close to $T_{d}+T_{\mathrm{RO}}=15.8$ µs, yielding a duty cycle $\eta_{\mathrm{cycle}}=0.95$. From Eq. 3, we expect $\eta_{\mathrm{SMPD}}$ to reach $0.74$.
The experimental $\eta_{\mathrm{SMPD}}$ is measured by counting the number of clicks detected over 1 second integration time, at the same time that a calibrated signal is sent to the SMPD input. The signal power is calibrated using the qubit as a detector (see [4] and Appendix B). The result is plotted as a function of input frequency $\omega$ in Fig. 2d. The maximum efficiency reaches $\eta_{\mathrm{SMPD}}=0.8$. This is even slightly higher than the predicted value, which implies in particular that the 4WM efficiency $\eta_{\mathrm{4WM}}$ is very close to $1$. Fitting $\eta_{\mathrm{SMPD}}(\omega)$ with a lorentzian, we obtain the detector bandwidth $\kappa_{d}$. Two curves are shown, obtained for two different values of $\kappa_{b}$. The maximum efficiency is unchanged, but the detector bandwidth is changed according to $\kappa_{b}$. This confirms that the SMPD bandwidth can be adjusted by tuning $\kappa_{b}$ via the Purcell filter frequency. We measure $\kappa_{d}/2\pi=315$ kHz for $\kappa_{b}/2\pi=170$ kHz, and $\kappa_{d}/2\pi=530$ kHz for $\kappa_{b}/2\pi=280$ kHz, thus confirming the approximate relation $\kappa_{d}\simeq 2\kappa_{b}$ (see Appendix A).
Dark counts are measured by recording photon-counting traces in the absence of any input signal (see Tab. 3, where the detector bandwidth was $\kappa_{d}/2\pi=170$ kHz). To obtain the various contributions to the dark counts, the traces are systematically measured with the pump off (yielding $\alpha_{q}$), with the pump on and detuned by $20$ MHz (yielding $\alpha_{q}+\alpha_{p}=\alpha_{\mathrm{err}}$), and with the pump on and tuned to the 4WM conditions (yielding $\alpha=\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$). At $10$ mK, we obtain $\alpha_{q}=8\mathrm{s}^{-1}$, $\alpha_{p}=2\mathrm{s}^{-1}$, and $\alpha_{\mathrm{th}}=21\mathrm{s}^{-1}$. Combining this with the measured efficiency, we obtain a power sensitivity $S=3\cdot 10^{-23}\mathrm{W}/\sqrt{\mathrm{Hz}}$, a factor of 3 better than the previous state of the art [4], which corresponds to an order of magnitude gain in measurement time.
Time traces are similarly measured by varying the cryostat temperature $T$. The values of $\alpha_{q}$ and $\alpha_{\mathrm{th}}$ are shown as a function of $T$ in Fig. 3. The qubit contribution increases with temperature because of the higher excited-state probability at equilibrium; it is well fitted by $\alpha_{q,0}+K_{q}p_{\mathrm{th},q}(T)$, with $\alpha_{q,0}=7\,\mathrm{s}^{-1}$ and $K_{q}=2.2\cdot 10^{4}\,\mathrm{s}^{-1}$. The value of $K_{q}$ is slightly larger than the predicted $(T_{d}/T_{1})(1/(T_{d}+T_{\mathrm{RO}}+T_{\mathrm{reset}}))$ (see Section IIIB), which is $1.4\cdot 10^{4}\,\mathrm{s}^{-1}$, possibly because of a reduction of $T_{1}$ at higher temperature. The thermal contribution also increases with temperature; it is well fitted by $\alpha_{\mathrm{th},0}+K_{\mathrm{th}}\bar{n}_{\mathrm{th},b}$, with $\alpha_{\mathrm{th},0}=31\mathrm{s}^{-1}$, and $K_{\mathrm{th}}=2\cdot 10^{5}\,\mathrm{s}^{-1}$. This value is in quantitative agreement with the expected $\eta_{\mathrm{SMPD}}\kappa_{d}/4=2.1\cdot 10^{5}\mathrm{s}^{-1}$, thus providing an independent confirmation of the efficiency estimate reported earlier.
The equilibrium qubit excitation (without active reset) $p_{\mathrm{th},q}$ at $10$ mK translates into an effective qubit temperature of $42$ mK. The thermal photon rate $\alpha_{\mathrm{th},0}$ measured at $10$ mK translates into a mean photon number per mode of $1.5\cdot 10^{-4}$ and thus to an effective temperature of $40$ mK. This is higher than the mean photon number expected from our line attenuation, which we estimate to be below $\sim 2\cdot 10^{-5}$. Our results therefore confirm the well-known difficulty of thermalizing microwave fields and modes to temperatures much lower than $\sim 40$ mK, possibly due to the difficulty of perfectly thermalizing microwave components. It is interesting to note that SMPD devices are ideally well suited to studying this phenomenon, which represents a major issue for superconducting quantum computers.
Temperature Pump state $\kappa_{d}/2\pi$ One-second click sequence Click rate
10 mK Off Not relevant [Uncaptioned image] 8/s
10 mK Detuned Not relevant [Uncaptioned image] 10/s
10 mK Tuned 170 kHz [Uncaptioned image] 31/s
50 mK Tuned 170 kHz [Uncaptioned image] 344/s
60 mK Tuned 170 kHz [Uncaptioned image] 621/s
90 mK Tuned 170 kHz [Uncaptioned image] 3614/s
Table 3: Photon-counting traces taken under different conditions. First line : the mixing-chamber temperature is 10 mK and the pump is off, yielding $\alpha_{q}$. Second line : the mixing-chamber temperature is 10 mK and the pump is switched on, but it is detuned by $20$ MHz from the 4WM condition, yielding $\alpha_{\mathrm{err}}=\alpha_{q}+\alpha_{p}$. Third line : the mixing-chamber temperature is 10 mK and the pump is on and in the 4WM condition, yielding $\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$. Three following lines : the mixing-chamber temperature is increased, and the count rate rises accordingly, as detailed in fig. 3.
Figure 3: Measured $\alpha_{\mathrm{th}}$ (blue full circles) and $\alpha_{q}$ (orange full circles) as a function of the cryostat’s temperature $T$, for a detection bandwidth $\kappa_{d}/2\pi=170$ kHz. The solid green line is a fit of $\alpha_{\mathrm{th}}$ to $\alpha_{\mathrm{th},0}+K_{\mathrm{th}}\bar{n}_{\mathrm{th},b}(T)$, yielding $\alpha_{\mathrm{th},0}=31\mathrm{s}^{-1}$ and $K_{\mathrm{th}}=2\cdot 10^{5}\mathrm{s}^{-1}$. The solid brown line is a fit of $\alpha_{q}$ to $\alpha_{q,0}+K_{q}p_{\mathrm{th},q}(T)$, yielding $\alpha_{q,0}=7\mathrm{s}^{-1}$ and $K_{q}=2.2\cdot 10^{4}\mathrm{s}^{-1}$. The dashed lines are the separated Bose-Einstein and constant components of the respective fits.
IV.4 Dependence on detection bandwidth
Figure 4: The blue full circles are the measured total dark counts, $\alpha_{\mathrm{err}}+\alpha_{\mathrm{th}}$, as a function of SMPD bandwidth $\kappa_{d}$. The dashed line is a linear fit, with slope $1.5\cdot 10^{-5}$, and intercept $11.6\mathrm{s}^{-1}$. The red triangles show the measured efficiency at resonance, $\eta_{\mathrm{SMPD}}$.
We finally study the SMPD figures of merit as a function of detection bandwidth, at the lowest temperature of $10$ mK. We measure the maximum value of $\eta_{\mathrm{SMPD}}(\omega)$ and $\alpha$ as a function of $\kappa_{d}$ (see Fig. 4). We observe that $\eta_{\mathrm{SMPD}}$ increases with $\kappa_{d}$, although the dependence is very weak as soon as the SMPD bandwidth is above $250$ kHz. This is expected because the measurements are performed with a monochromatic tone.The reduction in efficiency at the lowest value of $\kappa_{d}$ may be attributed to the contribution of internal buffer losses, as well as possible transient effects that effectively shorten the detection time $T_{d}$.
The false-positive rate $\alpha$, however, scales linearly with $\kappa_{d}$, as expected from Eq. 5. The fitted slope is $1.6\cdot 10^{-5}$, in agreement with the expected $\eta_{\mathrm{SMPD}}\bar{n}_{\mathrm{th},b}/4$ yielding a thermal mode occupation $\bar{n}_{\mathrm{th},b}=0.8\cdot 10^{-4}$, slightly lower than in the measurements of Fig. 3 and corresponding to a mode temperature of $38$ mK. The intercept at $\kappa_{d}=0$ is $11.6\,\mathrm{s}^{-1}$, consistent with the error rate $\alpha_{\mathrm{err}}$ measured with pump detuned. Overall, the observed linear dependence of $\alpha_{\mathrm{th}}$ on $\kappa_{d}$ proves that these counts are indeed due to thermal photons impinging on the SMPD input line. Not much can be inferred about their spectrum, apart from the fact that it appears to be flat on the scale of the detection bandwidth ($\sim 1$ MHz).
For fig. 4, no data have been taken above $\kappa_{d}/2\pi=$ 1 MHz. Indeed, above this value, high pump amplitudes are required to keep $C=1$, and our microwave setup was unable to send high enough power.
IV.5 Single-spin measurements
We confirm the high efficiency of the detector by measuring the microwave fluorescence from a single electronic spin. The setup closely follows refs [41] and [39], and is schematically depicted in Fig. 5. The electronic spin is an $\mathrm{Er}^{3+}$ ion in a $\mathrm{CaWO}_{4}$ matrix and is coupled to a superconducting resonator to enhance its radiative decay rate $\Gamma_{R}$. The spin is excited by a $\pi$ pulse with near-unit probability. It then relaxes radiatively by emitting a photon, which is either absorbed by the internal losses of the spin resonator, or emitted in the line, after which it may reach the SMPD detector or be absorbed by the losses between the two devices. The SMPD buffer $\omega_{b}$ is tuned to the spin and resonator frequency. The spin-to-click probability is therefore $\eta=\eta_{reso}\eta_{loss}\eta_{\mathrm{SMPD}}(\omega_{b})$, with $\eta_{reso}=0.6$ in this device (described in more detail elsewhere [39]).
The spin microwave fluorescence curve is shown in Fig. 5. This is well fitted by an exponential decay with time constant $\Gamma_{R}^{-1}=1.24$ ms on top of a constant background, due to the SMPD dark counts. From the integral below the fluorescence curve, we obtain $\eta=0.4$. This implies that the microwave power absorption between the spin device and the SMPD is $\eta_{loss}=0.85$, a very plausible value given that the line includes a superconducting cable, a circulator, and an infrared filter. This provides an independent confirmation of the high efficiency reached by this SMPD device and of its relevance for single-spin spectroscopy.
V Conclusion
We have demonstrated a microwave photon counter with tunable frequency (in the $7$ GHz range) and bandwidth (between $\sim 100$ kHz and $\sim 1$ MHz). The device has a maximum efficiency of $0.8$, and its noise can be reduced by lowering the bandwidth. At an input bandwidth of $200$ kHz, we demonstrate a record power sensitivity of $3\cdot 10^{-23}\mathrm{W}/\sqrt{\mathrm{Hz}}$. The device is well suited to detecting emitters with a narrow line, and it will find applications in single-spin magnetic resonance [39, 30] as well as in haloscope-based dark-matter searches [12].
Acknowledgements
We acknowledge technical support from P. Sénat, D. Duet, P.-F. Orfila, and S. Delprat, and we are grateful for fruitful discussions within the Quantronics group. We acknowledge support from the Agence Nationale de la Recherche (ANR) through the MIRESPIN (ANR-19-CE47-0011) project. We acknowledge support of the Région Ile-de-France through the DIM QUANTIP, from the AIDAS virtual joint laboratory, and from the France 2030 plan under Grant No. ANR-22-PETQ-0003. This project has received funding from the European Union Horizon 2020 research and innovation program under the OpenSuperQ100+ project, and from the European Research Council under Grant No. 101042315 (INGENIOUS). We acknowledge IARPA and Lincoln Labs for providing the Josephson traveling-wave parametric amplifier.
Figure 5: Microwave fluorescence of individual $\mathrm{Er}^{3+}$ electron spins in a $\mathrm{CaWO}_{4}$ host crystal matrix at $10$ mK, detected by the SMPD reported here operated with $\kappa_{d}/2\pi=500$ kHz (blue) and by a previous-generation SMPD operated with $\kappa_{d}/2\pi=400$ kHz (red). Top : scheme of the experiment. The $Er^{3+}$ spin is magnetically coupled to the microwave field $\vec{B}_{1}$ of an LC resonator and Zeeman tuned to resonance by a field $\vec{B}_{0}||$. The spin frequency is set to 7.7492 GHz in the latest version of the experiment, and 7.335 GHz for the ”old” version. The spin is excited by a $\pi$-pulse, and the photon it emits upon radiative relaxation is detected by the SMPD. Graph : blue full circles are the measured count rate following a $\pi$ pulse excitation of the spin, as a function of the time, with the SMPD presented herein. The solid blue line is an exponential fit, from which we extract the spin lifetime $\Gamma_{R}^{-1}=1.24$ ms and spin-to-click efficiency, $\eta=0.4$. Note that the noise floor $\alpha_{err}+\alpha_{th}=0.12\mathrm{ms}^{-1}$ is slightly higher than expected from Fig. 4, possibly due to a different wiring scheme in the cryostat. The red full circles and solid line are the equivalent data and fit for the same experiment performed with a previous-generation SMPD. One can observe a fourfold improvement of $\eta$ from the previous to the newest SMPD generation, yielding a measurement time faster by an order of magnitude.
[…]
Appendix A : Pump calibration
Figure 7: a) Four-wave mixing pattern : this color plot shows the measured qubit excited state as a function of pump frequency $\omega_{p}$ and signal frequency $\omega$ (from the input line). The qubit is excited only when the signal is close to $\omega_{b}$, and the pump verifies the 4WM condition. b) Fit from a two-coupled-cavity model. To fit the data, the value of the cooperativity $C$ is adjusted. Here, the value is $C=0.99$ (see appendix A). Other fit parameters are the buffer linewidth, $\nicefrac{{\kappa_{b}}}{{2\pi}}=120$ kHz, and the waste linewidth, $\nicefrac{{\kappa_{w}}}{{2\pi}}=1.75$ MHz. c) Same experiment as in a, but with the signal tone switched off. The higher excitation probabilities seen in subfigure a are no longer visible, demonstrating that we were indeed observing the actual 4WM pattern. Insets : corresponding pulse sequences. Purple pulses are sent through the pump/drive line, orange pulses through the signal/input line and green pulses through the waste/readout (RO) line. The pumping, signal, and readout pulses are respectively 15 µs, 10.5 µs and 800 ns long.
In this appendix we detail the tuning procedure for the pump parameters. The optimal pump frequency should satisfy the 4WM condition, $\omega_{\mathrm{4WM}}=\omega_{q}+\omega_{w}-\chi_{w}-\omega_{b}$. In the calibration procedure of the device, pump calibration comes after qubit characterization and readout calibration, which results are displayed in Fig. 6.
The pump amplitude is chosen by recording two-dimensional plots such as those shown in Fig. 7, where the qubit excited-state probability is recorded as a function of pump and signal frequency, $\omega_{p}$ and $\omega$, respectively, for various pump powers. Each plot is then fitted to the following formula [2] :
$|S_{w\leftarrow b}(\delta,\delta_{p})|^{2}=\frac{4C}{\bigg|1+C-4\frac{\delta(\delta+\delta_{p})}{\kappa_{b}\kappa_{w}}+2i\frac{\delta}{\kappa_{b}}+2i\frac{\delta+\delta_{p}}{\kappa_{w}}\bigg|^{2}}$ (6)
where $\delta=\omega-\omega_{b}$ is the signal-buffer detuning, and $\delta_{p}=\omega_{p}-\omega_{\mathrm{4WM}}$ the pump detuning from the 4WM condition.
If all detunings are zero, eq. 6 becomes :
$|S_{w\leftarrow b}(\delta=0,\delta_{p}=0)|^{2}=\frac{4C}{(1+C)^{2}}$ (7)
The maximum of eq. 7 is reached for $C=1$ and its value is 1.
The fit returns the value of $C$, and after a few iterations the condition $C=1$ is readily achieved.
Note that the full width at half maximum along the signal frequency axis of $|S_{w\leftarrow b}(\delta,\delta_{p})|^{2}$ is the detection bandwidth of the SMPD:
$\kappa_{d}=\sqrt{2}\sqrt{\sqrt{\kappa_{b}^{2}\kappa_{w}^{2}+\left(\frac{\kappa_{b}-\kappa_{w}}{2}\right)^{4}}-\left(\frac{\kappa_{b}-\kappa_{w}}{2}\right)^{2}}$ (8)
When $\kappa_{b}\ll\kappa_{w}$ (as is the case in this work), the formula can be approximated as
$\kappa_{d}\approx 2\kappa_{b}$ (9)
Figure 8: Detection bandwidth $\kappa_{d}$ as a function of buffer bandwidth $\kappa_{b}$. Full blue circles are experimental data and the blue solid line is a linear fit. The fit returns a slope of 1.9, which is close to 2, the value that is expected from eq. 9
Fig. 8 shows that eq. 9 is verified experimentally.
Appendix B : Efficiency measurement
Measuring the efficiency of the SMPD begins with estimating the power at the buffer input. This is achieved by using the qubit itself as a detector. Ramsey fringes are shifted by $\Delta_{q}$ and dephased by $\Gamma_{q}$ according to the equation:
$\Delta_{q}+i\Gamma_{q}=\frac{-4\chi_{b}|\epsilon|^{2}}{(\kappa_{b}+i\chi_{b})^{2}+4\delta^{2}},$ (10)
where $\epsilon$ is the complex amplitude of the intracavity field. The efficiency is then simply computed by dividing the number of clicks per second by the number of photons per second at the device input.
Appendix C : SMPD fabrication
We now describe the fabrication recipe of the SMPD presented in this paper.
C1 : Substrate preparation
The process starts with a two-inch, double-side-polished sapphire wafer. This wafer is dipped in a piranha solution made with 2/3 sulfuric acid and 1/3 hydrogen peroxyde (by volume) and left for 20 minutes. Then, it is rinsed with water and dried with a stream of nitrogen. After that, the wafer is loaded into an RF reactive sputterer together with a tantalum target. The targeted thickness of the tantalum sputtering is between 80 and 90 nm. The sputtering process shall be performed while the wafer is heated up to 600°C to obtain alpha phase tantalum.
The wafer is then prepared for dicing. To this end, UV III is spin-coated over the wafer for 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s, followed by a postbake at 120°C for 3 minutes.
C2 : Circuit etching
After dicing, the protective UV III layer is removed applying isopropanol (IPA) from a wash bottle the chosen chip for 15 seconds, then again for 5 seconds. For further cleaning, the chip is immersed in acetone under strong ultrasonic agitation for 5 minutes, then in another ultrasonic bath of IPA for 3 minutes. In the following, it is then dried with a stream of nitrogen and immersed in a piranha solution for 20 minutes. The chip is then rinsed with water, dehydrated with IPA, dried with a stream of nitrogen, and further dehydrated on a hotplate at 120°C for 5 minutes. Next, AZ1518 is spin-coated for 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. The postbake is performed on a hotplate for 2 minutes at 115°C. The circuit is then patterned using a UV LASER direct writer. A supplementary exposure is needed at the edge bead. The development is performed in pure MF319 for 60 seconds and stopped with water. The chip is dried with a stream of nitrogen and hard baked on a hotplate for 2 minutes at 115°C. Finally, the circuit is etched by dipping in Transene 111 (1/3 nitric acid, 1/3 hydrofluoric acid, and 1/3 water) for exactly 20 seconds. There is no need for agitation during etching. Rinsing is performed in a first bath of water for 20 seconds with agitation and then in a second bath of water for transport between benches.
C3 : Chip cleaning and oxide removal
After the etching step, the chip is re-cleaned before proceeding to the e-beam related steps. Recently, we have added an oxide-removal step to the process which has not been applied to the sample presented here. First, the sample is immersed in a bath of acetone for 5 minutes to remove AZ1518, then in another bath of IPA for 3 minutes, and finally in piranha solution for 2 minutes. The chip is rinsed with water, dehydrated with IPA and dried with a stream of nitrogen.
In the recently added oxidation step, the chip is immersed for 20 minutes in buffered oxide etchant (BOE), rinsed in a first bath of water for 20 seconds with agitation, then in a second bath of water for transport between benches, dehydrated with IPA, dried with a stream of nitrogen, and further dehydrated on a hotplate during 5 minutes at 120°C.
C4 : Preparation of e-beam bilayer
A first layer of MAA EL-6 is spin-coated for 60 seconds at 4000 rpm with an acceleration of 1000 rpm/s. The postbake is performed on a hotplate for 5 minutes at 172°C. This is followed by spin coating of a second layer of PMMA 950k A6 for 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. The postbake, on a hotplate at 175°C, lasts 15 minutes.
The substrate being made of sapphire, there is no way for electrons to escape through it during lithography. Therefore, the bilayer must be metallized to prevent charge effects. For this reason, a 7-nm thick layer of aluminum is evaporated over the bilayer at a rate of 0.2 nm/s. There is no need for a good vacuum for this evaporation step ; a pressure of about $10^{-6}$ hPa is enough.
C5 : E-beam lithography
Given the need for the simultaneous presence of Josephson junctions of very different sizes across the device, the Manhattan process has been chosen. Due to the absence of a 100-kV electron beam lithography (EBL) system in our laboratory, we developed an adapted version of this process using a bilayer with a reversed aspect ratio (a layer of MAA about 170 nm thick beneath a layer of PMMA about 580 nm thick). We use the following parameters with our Raith e-LINE system :
  • Magnification : $\times 1000$.
  • Base dose : $250\mu C/cm^{2}$.
  • Relative dose factor : 1.5 for ”big” junctions’ main pattern (about 1 µm edge size), 1.7 for ”small” junctions’ main pattern (about 200 nm edge size or less), 0.3 for undercut boxes.
  • Voltage : 25 kV.
  • Aperture : 10 µm.
  • Current : About 20 pA.
  • Writefield : $100\times 100\mu m^{2}$.
C6 : Aluminum etching, resist development, and descum
Before developing the e-beam resist, the metallization layer must be removed. Aluminum etching is performed by dipping the chip in a 10 g/l potassium hydroxyde (KOH) solution in water for 30 seconds. It is then rinsed by for 3 seconds in water, followed by 3 seconds in IPA. The development itself begins immediately after : The chip is immersed for 45 seconds in a solution made of 1/4 methyl isobutyl ketone (MIBK) and 3/4 IPA (by volume). Development is stopped by dipping in IPA for 15 seconds. The chip and the bilayer are dried first with a stream of nitrogen, then on a hotplate at 70°C for 1 minute. Finally, the resist mask is cleaned in an oxygen plasma asher for 10 seconds, at a pressure of 0.2 hPa and an ionizing power of 75 W.
C7 : Josephson junctions evaporation
For evaporating Josephson junctions, a good vacuum is required (at least below $5\cdot 10^{-7}$ hPa). Given that we are using the Manhattan process, the evaporator must be equipped with a planetary rotation. The parameters given here apply to one of our Plassys MEB 550S systems.
First, the surface is prepared by argon-ion milling with an incidence angle of 0°. The ion gun is tuned with the filament at 500 V and 25 mA, and the acceleration grid at 100 V. The shutter is removed from the ion-beam trajectory twice for 5 seconds each, with 1 minute of rest between exposures. Then comes the evaporation of the first layer of aluminum, performed at an angle of 30° and a rate of 1 nm/s. A thickness of 29 nm is set on the evaporator to obtain an actual thickness of 25 nm. This first layer is oxidized by a mixture of oxygen and argon at 10 hPa for 5 minutes. The sample holder is rotated around the planetary axis by 90° in order to evaporate along the orthogonal direction of the Manhattan pattern. Finally, the second evaporation of aluminum is performed with the same parameters (30° angle and 1 nm/s rate), except that a thickness of 69 nm is is set to get an actual thickness of 60 nm.
C8 : Final lift off
The lift off is performed in two successive baths of acetone at 60°C. Blowing with a Pasteur pipette helps for removal. The sample is then brought back to room temperature with a non-heated bath of acetone. It is then dehydrated in IPA and finally dried by dinitrogen blow.
Appendix D : wiring diagram
The wiring scheme used for SMPD characterization is presented in fig. 9. Lines ”In A1” and ”Out 2” are used to probe the buffer resonator in reflection (to determine $\omega_{b}$ and $\kappa_{b}$), and also to calibrate the detector efficiency by sending pulses containing a well-controlled number of photons. ”In B1” and ”In B4” are flux bias lines for controlling the SQUID inductances of the buffer and its Purcell resonator, respectively. ”In A2” and ”Out 3” are used for probing the waste resonator in reflection (to determine $\omega_{w}$ and $\kappa_{w}$), and also to perform dispersive readout of the qubit through the waste resonator. ”In B2” is used to pump the travelling-wave parametric amplifier (TWPA) on the ”Out 3” line. ”In B3” is the qubit drive line.
Figure 9: Wiring used for the characterization run reported herein.